Observations are taken on sales of a certain mountain bike in 24 sporting goods stores. The regression model was Y = total sales (thousands of dollars), X1 = display floor space (square meters), X2 = competitors' advertising expenditures (thousands of dollars), X3 = advertised price (dollars per unit).

Predictor Coefficient SE tcalc p-value

Find the T calc and P-Value of the following;

Intercept 1,265.1 343.6
FloorSpace 11.455 1.54
Competing Ads -6.453 3.908
Price -0.14956 0.08880

(b-1) What is the critical value of Student's t in Appendix D for a two-tailed test at α = .01? (Round your answer to 3 decimal places.)

t-value = ______
Answer #1

tCritical value for a = 0.01 , d.f= n-2 = 22 and two tailed test

tCritical = t​​​​​​a/2, n-2  = t​​​​​​0.005,22

tCritical = 2.819

t-value = 2.819